OpenStudy (anonymous):

Find the probability of rolling a sum of seven with a single toss of two fair six-sided dice. Select one: a. 0 b. 1/7 c. 1/6 d. 1/9

OpenStudy (anonymous):

D, or B

OpenStudy (anonymous):

I think its D. for sure?

jimthompson5910 (jim_thompson5910):

a chart like this may help https://bestcase.files.wordpress.com/2011/01/dicediagram.jpg

jimthompson5910 (jim_thompson5910):

count how many times the sum 7 comes up divide that count by 36 (since there are 6*6 = 36 different ways to roll 2 dice)

OpenStudy (anonymous):

Okay thank you :)

OpenStudy (anonymous):

c.

OpenStudy (anonymous):

1,6 2,5 3,4 6,1 5,2 4,3 6/36 = 1/6

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Your welcome and thank you for the medal.